# Pathwise uniform convergence of a full discretization for a three-dimensional stochastic Allen-Cahn equation with multiplicative noise

### v4

@@ 3.2 A stability estimate for a discrete stochastic convolution
$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}c\sum_{j=0}^{J-1}\mathbb{E}\lVert{\delta_{j}}\rVert_{L^{q}}^{p}+c\sum_{j=0}^{J-1}\tau^{p/2-1}\int_{t_{j}}^{t_{j+1}}\lVert{\Psi(t)}\rVert_{L^{p}(\Omega;\gamma(H,L^{q}))}^{p}\,\mathrm{d}t$
@@
$\stackrel{{\scriptstyle\mathrm{(iv)}}}{{\leqslant}}c\tau^{p/2}\sum_{j=0}^{J-1}\mathbb{E}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}+c\tau^{p/2-1}\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},$
@@
where in (i) we use the definition $M_{j}:=\delta_{j}+\int_{t_{j}}^{t_{j+1}}\Psi(t)\,\mathrm{d}W_{H}(t)$, in (ii) we invoke Lemma 3.3, in (iii) we apply Hölder’s inequality, and in (iv) we utilize the definition $\delta_{j}:=(I-\tau\Delta_{h})^{-1}Z_{j}-Z_{j}$, along with Eq. 3.3 with $\alpha=0$ and $\beta=1$. The desired inequality Eq. 3.13 then follows by combining the above estimates.
Step 5. Combining Eqs. 3.10, 3.11, 3.12 and 3.13 yields
$\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}$ $\leqslant\varepsilon\mathbb{E}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}+c\tau\sum_{j=0}^{J-1}\mathbb{E}\Big{(}\lVert{Z_{j}}\rVert_{L^{q}}^{p}+\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\Big{)}$
$\qquad{}+c(1+\varepsilon^{1-p})\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},\quad\forall\varepsilon\in(0,1).$
Invoking the initial condition $Z_{0}=0$ and exploiting the continuous embedding of $\dot{H}_{h}^{1,q}$ into $L^{q}$, we deduce
@@
$\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\right]\leqslant c\tau\sum_{j=1}^{J-1}\mathbb{E}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}+c\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}.$
@@
$\tau\sum_{j=1}^{J-1}\mathbb{E}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\leqslant c\lVert{P_{h}g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}.$
@@
$\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}\leqslant c\lVert{P_{h}g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}.$
Aggregating these estimates and invoking Lemma 3.1, we conclude
$\mathbb{E}\Big{[}\max_{1\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},$
which implies the stability estimate (3.8). This completes the proof. ∎
@@
In this subsection, we always assume that the conditions of Theorem 2.1 are satisfied. Let $y_{h}$ be the solution of the following spatial semi-discretization:
$\begin{cases}{}\mathrm{d}y_{h}(t)=(\Delta_{h}y+y_{h}-P_{h}y_{h}^{3})(t)\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad t>0,\\ y_{h}(0)=P_{h}v.\end{cases}$ (3.14)
It is standard that
$y_{h}(t)$ $=e^{t\Delta_{h}}P_{h}v+\int_{0}^{t}e^{(t-s)\Delta_{h}}(y_{h}-P_{h}y_{h}^{3})(s)\,\mathrm{d}s+{}$ (3.15)
$\qquad\int_{0}^{t}e^{(t-s)\Delta_{h}}P_{h}F(y_{h}(s))\,\mathrm{d}W_{h}(s),\quad t\in[0,T].$
Moreover, Li and Zhou [22, Theorem 3.3] have established the error estimate
@@
for all $p\in(2,\infty)$, $q\in[2,\infty)$ and $\varepsilon>0$. Hence, to conclude the proof of Theorem 2.1, it remains to prove
@@
which is the main task of the rest of this subsection.
To this end, we first note that the solution $y$ possesses the following regularity, as indicated by [22, Equation (4.42)]:
$y\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0.$
Using this regularity result, along with the stability property of $P_{h}$ as stated in Lemma 3.1, the standard approximation properties of $P_{h}$, the error estimate Eq. 3.16, and the inverse estimate
$\lVert{v_{h}}\rVert_{\dot{H}_{h}^{1,q}}\leqslant ch^{-1}\lVert{v_{h}}\rVert_{L^{q}},\quad\forall v_{h}\in X_{h},\,\forall q\in[2,\infty),$
we deduce the following regularity assertion for $y_{h}$:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))}+\lVert{P_{h}y_{h}^{3}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))}$ (3.18)
@@
Nevertheless, it is important to acknowledge that the norm of $y_{h}$ in the space $L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))$ depends on the regularity parameters of the spatial triangulation $\mathcal{K}_{h}$.
@@
    In Subsection 3.3.1 we derive the following estimates:
    $\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}(t)-y_{h}(t_{j})}\rVert_{L^{q}}^{p}\,\mathrm{d}t\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty),$ (3.19)
    $\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{L^{q}}^{p}\,\mathrm{d}t\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty),$ (3.20)
    $\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{\dot{H}_{h}^{1,q}}^{p}\,\mathrm{d}t\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty),$ (3.21)
    $\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\rVert_{L^{q}}^{p}\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty).$ (3.22)
@@
    In Subsection 3.3.2, we derive the error estimate
    $\max_{0\leqslant j\leqslant J}\lVert{y_{h}(t_{j})-Y_{j}}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\tau^{1/2},\quad\forall p\in[2,\infty).$ (3.23)
@@
    In Subsection 3.3.3, we establish the stability estimate
    $\hskip-25.00003pt\tau\sum_{j=1}^{J}\mathbb{E}\lVert{Y_{j}}\rVert_{L^{q}}^{p}\text{ is uniformly bounded w.r.t.~{}$h$ and $\tau$ for all $p,q\in[2,\infty)$}.$ (3.24)
@@
    In Subsection 3.3.4, we establish the error estimate
    $\tau\sum_{j=0}^{J}\mathbb{E}\lVert{Y_{j}-y_{h}(t_{j})}\rVert_{L^{q}}^{p}\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty).$ (3.25)
  5. 5.
    Finally, in Subsection 3.3.5, by utilizing Eqs. 3.18, 3.19, 3.20, 3.21, 3.22, 3.23, 3.24 and 3.25, along with the discrete stochastic maximal $L^{p}$-regularity estimate from Proposition 3.1 and the stability estimate from Proposition 3.2, we conclude the proof for the desired error estimate (3.17).
3.3.1 Proof of the estimates Eqs. 3.19, 3.20, 3.21 and 3.22
In light of the regularity assertion Eq. 3.18 and the inequalities Eqs. 3.1 and 3.2, the derivation of the estimates Eqs. 3.19, 3.20, 3.21 and 3.22 proceeds in a manner that is both direct and routine. Consequently, to maintain brevity, we shall hereinafter provide only a concise proof specifically for the case of Eq. 3.21.
To this end, we proceed as follows. Fix any $0\leqslant j<J$ and $t\in[t_{j},t_{j+1}]$. From Eq. 3.15 we deduce the decomposition
$y_{h}(t_{j+1})-y_{h}(t)=I_{0}+I_{1}+I_{2}+I_{3}+I_{4},$
where each term is defined as follows:
$\hskip-125.00018ptI_{0}:=\Big{(}e^{t_{j+1}\Delta_{h}}-e^{t\Delta_{h}}\Big{)}P_{h}v,$
$\hskip-45.00006ptI_{1}:=\int_{t}^{t_{j+1}}e^{(t_{j+1}-s)\Delta_{h}}\big{(}y_{h}(s)-P_{h}y_{h}^{3}(s)\big{)}\,\mathrm{d}s,$
$I_{2}:=\int_{0}^{t}\Big{(}e^{(t_{j+1}-s)\Delta_{h}}-e^{(t-s)\Delta_{h}}\Big{)}\big{(}y_{h}(s)-P_{h}y_{h}^{3}(s)\big{)}\,\mathrm{d}s,$
$\hskip-50.00008ptI_{3}:=\int_{t}^{t_{j+1}}e^{(t_{j+1}-s)\Delta_{h}}P_{h}F(y_{h}(s))\,\mathrm{d}W_{H}(s),$
$\hskip-5.0ptI_{4}:=\int_{0}^{t}\Big{(}e^{(t_{j+1}-s)\Delta_{h}}-e^{(t-s)\Delta_{h}}\Big{)}P_{h}F(y_{h}(s))\,\mathrm{d}W_{H}(s).$
Using the inequalities Eq. 3.1 and Eq. 3.2, we obtain the following bounds for the terms in the decomposition:
$\lVert{I_{0}}\rVert_{L^{p}(\Omega;\dot{H}_{h}^{1,q})}$ $\leqslant c\tau^{1/2}\lVert{P_{h}v}\rVert_{\dot{H}_{h}^{2,q}},$
$\lVert{I_{1}}\rVert_{L^{p}(\Omega;\dot{H}_{h}^{1,q})}$ $\leqslant c\tau\lVert{y_{h}-P_{h}y_{h}^{3}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))},$
$\lVert{I_{2}}\rVert_{L^{p}(\Omega;\dot{H}_{h}^{1,q})}$ $\leqslant c\tau^{1/2}\lVert{y_{h}-P_{h}y_{h}^{3}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))},$
$\lVert{I_{3}}\rVert_{L^{p}(\Omega;\dot{H}_{h}^{1,q})}$ $\leqslant c\tau^{1/2}\lVert{P_{h}F(y_{h})}\rVert_{L^{p}(\Omega;C([0,T];\gamma(H,\dot{H}_{h}^{1,q}))},$
$\lVert{I_{4}}\rVert_{L^{p}(\Omega;\dot{H}_{h}^{1,q})}$ $\leqslant c\tau^{1/2}\lVert{P_{h}F(y_{h})}\rVert_{L^{p}(\Omega;C([0,T];\gamma(H,\dot{H}_{h}^{1,q}))}.$
Drawing upon the stability property of the projection $P_{h}$ as established in Lemma 3.1, the initial condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, the growth estimate Eq. 3.6, and the regularity assertion Eq. 3.18, we demonstrate that the sum of the norms
$\lVert{P_{h}v}\rVert_{\dot{H}_{h}^{2,q}}+\lVert{y_{h}-P_{h}y_{h}^{3}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))}+\lVert{P_{h}F(y_{h})}\rVert_{L^{p}(\Omega;C([0,T];\gamma(H,\dot{H}_{h}^{1,q})))}$
remains uniformly bounded with respect to the mesh size $h$. Consequently, we combine these results to conclude that
$\lVert{y_{h}(t_{j+1})-y_{h}(t)}\rVert_{L^{p}(\Omega;\dot{H}_{h}^{1,q})}\leqslant c\tau^{1/2}$
for all $t\in[t_{j},t_{j+1}]$ with $0\leqslant j<J$. Subsequently, it follows that
$\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{\dot{H}_{h}^{1,q}}^{p}\,\mathrm{d}t\leqslant c\sum_{j=0}^{J-1}\tau^{1+p/2}\leqslant c\tau^{p/2},$
thereby establishing the desired estimate Eq. 3.21.
@@
Let $e_{j}:=Y_{j}-y_{h}(t_{j})$ for each $0\leqslant j\leqslant J$. Given Eq. 3.14, we have
$y_{h}(t_{j+1})-y_{h}(t_{j})$ $=\int_{t_{j}}^{t_{j+1}}\Delta_{h}y_{h}(t)+y_{h}(t)-P_{h}y_{h}^{3}(t)\,\mathrm{d}t+P_{h}\int_{t_{j}}^{t_{j+1}}F(y_{h}(t))\,\mathrm{d}W_{H}(t),$
from which, by Eq. 2.2, we deduce that
$e_{j+1}-e_{j}$ $=\int_{t_{j}}^{t_{j+1}}\Delta_{h}\big{(}Y_{j+1}-y_{h}(t)\big{)}+Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\,\mathrm{d}t$ (3.26)
@@
The remainder of the proof is organized into the following three steps.
Step 1. Let us prove
$\mathbb{E}\lVert{e_{j+1}}\rVert_{L^{2}}^{2m}\leqslant(1+c\tau)\mathbb{E}\big{\lVert}{e_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\big{\rVert}_{L^{2}}^{2m}$ (3.27)
$\qquad{}+c\tau\mathbb{E}\lVert{e_{j}}\rVert_{L^{2}}^{2m}+c\mathbb{E}\big{(}A_{j,m}+B_{j,m}+C_{j,m}\big{)},\quad\forall m\in\mathbb{N}_{>0},\,\,\forall 0\leqslant j<J,$
@@
$A_{j,m}$ $:=\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j})}\rVert_{L^{2}}^{2m}\,\mathrm{d}t,$
$B_{j,m}$ $:=\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{\dot{H}_{h}^{1,2}}^{2m}\,\mathrm{d}t,$
$C_{j,m}$ $:=\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\rVert_{L^{2}}^{2m}\,\mathrm{d}t.$
Fix any $m\in\mathbb{N}_{>0}$ and $0\leqslant j<J$. Consider the function
$g(\theta)$ $:=e_{j}+\theta\int_{t_{j}}^{t_{j+1}}\Big{[}\Delta_{h}\big{(}Y_{j+1}-y_{h}(t)\big{)}+Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\Big{]}\,\mathrm{d}t$

### v5

@@ 4.2 A stability estimate for a discrete stochastic convolution
$\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{M_{j}}\rVert_{L^{q}}^{p}\bigg{]}$ $\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{\delta_{j}}\rVert_{L^{q}}^{p}\bigg{]}+c\tau^{p/2-1}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{\Psi(t)}\rVert_{L^{p}(\Omega;\gamma(H,L^{q}))}^{p}\,\mathrm{d}t$
@@
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\tau^{p/2}\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\bigg{]}+c\tau^{p/2-1}\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},$
@@
where in the first inequality (i), we invoke the definition $M_{j}:=\delta_{j}+\int_{t_{j}}^{t_{j+1}}\Psi(t)\,\mathrm{d}W_{H}(t)$, Lemma 4.3, and Hölder’s inequality; and in the second inequality (ii), we make use of the definition $\delta_{j}:=(I-\tau\Delta_{h})^{-1}Z_{j}-Z_{j}$, along with the resolvent estimate from Lemma 4.3 with $\alpha=0$ and $\beta=1$. By combining these estimates, we obtain
@@
$I_{3}+I_{4}\leqslant c\mathbb{E}\bigg{[}\tau\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\bigg{]}+c\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}.$ (4.13)
Step 5. Combining Eqs. 4.10, 4.11, 4.12 and 4.13, and noting that $Z_{0}=0$, we deduce that
$\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\right]\leqslant c\mathbb{E}\bigg{[}\tau\sum_{j=1}^{J-1}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\bigg{]}+c\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}.$
@@
$\mathbb{E}\bigg{[}\tau\sum_{j=1}^{J-1}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\bigg{]}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))}^{p}.$
@@
$\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))}.$
Aggregating these estimates, we conclude the desired stability estimate (4.8). This completes the proof. ∎
Remark 4.1.
Considering the maximal inequality established in [40, Theorem 1.1],
$\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\left\lVert\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s)\right\rVert_{L^{q}}^{p}\bigg{]}\leqslant c\mathbb{E}\left(\int_{0}^{T}\lVert g(t)\rVert_{\gamma(H,L^{q})}^{2}\,\mathrm{d}t\right)^{p/2},$
for all $g\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))$ with $p,q\in[2,\infty)$, the stability estimate provided in Proposition 4.2 is suboptimal. While Proposition 4.2 is adequate for the purposes of the current investigation, it is an interesting question to derive a sharper stability estimate than the one given by (4.8).
@@
In this subsection, we always assume that the conditions of Theorem 3.1 are satisfied. Let $y_{h}$ be the strong solution of the following spatial semi-discretization:
$\begin{cases}\mathrm{d}y_{h}(t)=(\Delta_{h}y_{h}+y_{h}-P_{h}y_{h}^{3})(t)\,\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),&t\in[0,T],\\ y_{h}(0)=P_{h}v.\end{cases}$ (4.14)
Li and Zhou [28, Theorem 3.3] have established the error estimate
@@
for all $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$. Therefore, to complete the proof of Theorem 3.1, it remains to prove
@@
which is the primary focus of the remainder of this subsection.
To this end, we first note that by the regularity property of the mild solution $y$ as stated in Proposition 2.1, the stability property of $P_{h}$ as stated in Lemma 4.1, the standard approximation properties of $P_{h}$, the error estimate (4.15), the growth estimate of $F$ in (4.6), and the inverse estimate, we deduce the following stability assertion for $y_{h}$:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{2-\epsilon,q}))}+\lVert{P_{h}y_{h}^{3}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))}+\lVert{P_{h}F(y_{h})}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))}$ (4.17)
@@
With this stability assertion and the inequalities (4.1) and (4.2), a routine calculation yields the following estimates:
$\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j})}\rVert_{L^{q}}^{p}\,\mathrm{d}t\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty),$ (4.18)
$\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{L^{q}}^{p}\,\mathrm{d}t\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty),$ (4.19)
$\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{\dot{H}_{h}^{1/2+1/p,q}}^{p}\,\mathrm{d}t\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty),$ (4.20)
$\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\rVert_{L^{q}}^{p}\,\mathrm{d}t\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty).$ (4.21)
@@
    In Subsection 4.3.1, we derive the error estimate
    $\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\lVert{y_{h}(t_{j})-Y_{j}}\rVert_{L^{2}}^{p}\biggr{]}\leqslant c\tau^{p/2},\quad\forall p\in[2,\infty).$ (4.22)
@@
    In Subsection 4.3.2, we establish the following stability property:
    $\mathbb{E}\bigg{[}\tau\sum_{j=1}^{J}\lVert{Y_{j}}\rVert_{L^{q}}^{p}\biggr{]}\text{ is uniformly bounded w.r.t.~{}$h$ and $\tau$ for all $p,q\in(2,\infty)$}.$ (4.23)
@@
    In Subsection 4.3.3, we establish the error estimate
    $\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\lVert{Y_{j}-y_{h}(t_{j})}\rVert_{L^{q}}^{p}\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty).$ (4.24)
@@
    Finally, in Subsection 4.3.4, by utilizing Eqs. 4.17, 4.18, 4.19, 4.20, 4.21, 4.22, 4.23 and 4.24, along with the discrete deterministic maximal $L^{p}$-regularity estimate from Lemma 4.5, the discrete stochastic maximal $L^{p}$-regularity estimate from Proposition 4.1, and the stability estimate from Proposition 4.2, we conclude the proof for the desired error estimate (4.16).
@@
For each $0\leqslant j\leqslant J$, let $e_{j}:=Y_{j}-y_{h}(t_{j})$. It is clear that
$e_{j+1}-e_{j}$ $=\int_{t_{j}}^{t_{j+1}}\Delta_{h}\big{[}Y_{j+1}-y_{h}(t)\big{]}+Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\,\mathrm{d}t$ (4.25)
@@
We introduce the decomposition $e_{j}=\xi_{j}+\eta_{j}$ for each $0\leqslant j\leqslant J$, defining the sequence $(\xi_{j})_{j=0}^{J}$ as follows:
$\begin{cases}\xi_{j+1}-\xi_{j}=\tau\Delta_{h}\xi_{j+1}+\int_{t_{j}}^{t_{j+1}}\Delta_{h}\big{[}y_{h}(t_{j+1})-y_{h}(t)\big{]}\,\mathrm{d}t,\quad 0\leqslant j<J,\\ \xi_{0}=0.\end{cases}$
It can be readily verified that
$\eta_{j+1}-\eta_{j}$ $=\int_{t_{j}}^{t_{j+1}}\left[\Delta_{h}\eta_{j+1}+Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\right]\,\mathrm{d}t$ (4.26)
Moreover, by Lemma 4.5 and Eq. 4.19 we obtain
$\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\lVert{\xi_{j}}\rVert_{L^{q}}^{p}\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty).$ (4.27)
The remainder of the proof is organized into the following four steps.
Step 1. Let us prove that, for any $m\in\mathbb{N}_{>0}$ and $0\leqslant j<J$,
$\mathbb{E}\lVert{\eta_{j+1}}\rVert_{L^{2}}^{2m}\leqslant(1+c\tau)\mathbb{E}\Big{\lVert}{\eta_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{2}}^{2m}$ (4.28)
$\quad{}+c\tau\mathbb{E}\lVert{\eta_{j}}\rVert_{L^{2}}^{2m}+c\tau\mathbb{E}\lVert{\xi_{j}}\rVert_{L^{2}}^{2m}+c\mathbb{E}\big{[}A_{j,m}+B_{j,m}\big{]},$
@@
$A_{j,m}$ $:=\tau\langle{\xi_{j+1}^{2},\,y_{h}^{2}(t_{j+1})+Y_{j+1}^{2}}\rangle^{m},$
$B_{j,m}$ $:=\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j})}\rVert_{L^{2}}^{2m}+\lVert{y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\rVert_{L^{2}}^{2m}\,\mathrm{d}t.$
To this end, fix any $m\in\mathbb{N}_{>0}$ and $0\leqslant j<J$. Consider the function
$g(\theta)$ $:=\eta_{j}+\theta\int_{t_{j}}^{t_{j+1}}\Big{[}\Delta_{h}\eta_{j+1}+Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\Big{]}\,\mathrm{d}t$